OpenStudy (anonymous):

The has tank in hana's car holds 15 1/2 gallons. if hana used 3/4 of a tank of gas when she drove across Arizona , how many gallons of gas did she use on that trip

OpenStudy (anonymous):

The gas

OpenStudy (anonymous):

Convert to decimal, multiply, covert back to fraction.

OpenStudy (anonymous):

Is it subtraction the problem

OpenStudy (anonymous):

Leonhart

OpenStudy (anonymous):

No. Your simply trying to figure out what 3/4 is of 15 1/2. Convert to decimal. 15.50 0.75 Multiply. 15.5 * .75 Convert to fraction.

OpenStudy (anonymous):

You would subtract if it asked how much she had left

OpenStudy (anonymous):

So 11 1/4

OpenStudy (anonymous):

15.5 * .75 = 11.625 11 1/4 would = 11.25

OpenStudy (anonymous):

Can u do that in fractions not decmials

OpenStudy (anonymous):

0.625 = 5/8. Your answer is 11 5/8 :)

OpenStudy (anonymous):

So u would do 15 1/2 equals 31/2 then muiltply that to 3/4 and then get 93/8 then get 11 4/8 simplfy equals 11 1/4 tell me what i did wrong

OpenStudy (anonymous):

First, convert the fractions (15 1/2 & 3/4) to decimals. Then multiply the two together. (15.5 * .75) Then convert the decimal to a fraction. (11.625 = 11 5/8)

OpenStudy (anonymous):

If you don't want to convert to decimals and back into a fraction you can simply multiply the fractions. Whatever works best for you. Multiplying the fractions: Convert 15 1/2 to a improper fraction. = 31/2 Multiply 31/2 by 3/4. = 93/8 Covert/simplify that into a mixed number. = 11 5/8